Fermi National Laboratory
 

Accelerator Update for February 10, 2006 - February 13, 2006

 

Friday February 10

The day shift began with Operations monitoring store 4634 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the store, and with NuMI and MiniBooNE taking beam.

At 10:05 AM, Operations turned off the proton beam so that experts could begin work on the Main Injector (MI), a Linac RF station, and in Booster.

A Linac expert completed his work on LRF3 and Booster experts completed their work on SEXTL and DOGL3 at 10:33 AM. MiniBooNE resumed taking beam.

MI experts completed their work on the RF system at 11:05 AM. NuMI resumed taking beam.

Operations established beam to Meson MTest experiment at 11:47 AM.

Operations began preparing the Recycler to stash antiprotons at 1:01 PM. The stash ended and MTest resumed taking beam at 1:59 PM.

Operations terminated store 4634 at 3:57 PM. Upon termination, the Tevatron (TeV) suffered a sector F4 quench.

Operations established beam to the Meson MCenter experiment at 4:28 PM.

Operations prepared D0 and CDF for collision hall accesses at 4:33 PM.

CDF came out off access at 6:04 PM and D0 came out around 6:35 PM.

The Recycler began preparing for a stash at 6:56 PM.

Operations put the TeV through a dry squeeze at 7:10 PM.

The stash ended at 7:14 PM.

Pbar resumed stacking at 7:30 PM.

Operation put the TeV into shot setup at 7:46 PM.

At 9:59 PM, the TeV suffered a sector E4 quench due to a large amount of DC beam during the first transfer of antiprotons.

Pbar resumed stacking at 10:02 PM.

Operations put the TeV into shot setup at 11:53 PM.

Saturday February 11

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash and stack, and with NuMI and MiniBooNE taking beam.

Operations established store 4638 at 1:46 AM with an initial luminosity of 140.6E30.

The Recycler began preparing to stash at 1:55 AM. The stash ended at 2:22 AM. Pbar resumed stacking and MCenter resumed taking beam.

The Recycler began preparing to stash at 6:11 AM. The stash ended at 6:43 AM. Pbar resumed stacking and MCenter resumed taking beam.

At 8:40 AM, a Linac RF station (LRF2) tripped off. Operators investigated and then called in an expert.

Operations reported that the expert had repaired LRF2 and that Linac resumed sending beam at 11:57 AM.

At 12:02 PM, LRF2 tripped off for the first of six times in forty minutes. It reset each time.

The Recycler began preparing to stash at 1:45 PM. The stash ended at 2:27 PM. Pbar resumed stacking and MCenter resumed taking beam.

The Recycler began preparing to stash at 5:30 PM. The stash ended at 6:04PM. Pbar, MTest, and MCenter resumed taking beam at 6:13 PM.

The MTest experiment finished using the beam at 8:22 PM.

The Recycler began preparing to stash at 9:52 PM. The stash ended at 10:15 PM. Pbar resumed stacking and MCenter resumed taking beam.

At 11:48 PM, the Recycler suffered a significant loss of the stash. Investigation found that a quadrupole compensation circuit had come into alarm on a tracking error.

Sunday February 12

The midnight shift began with Operations monitoring store 4638 and the Recycler stash, with Pbar stacking, and with NuMI, MiniBooNE, and MCenter taking beam.

At 12:11 AM, the Recycler had another beam loss. Investigation discovered a horizontal focusing error. Recycler and MI experts were notified.

The Recycler began preparing to stash at 1:54 AM. The stash ended at 2:29 AM. Pbar resumed stacking and MCenter resumed taking beam at 2:42 AM.

The Recycler began preparing to stash at 6:47 AM. The stash ended at 7:12 AM. Pbar resumed stacking and MCenter resumed taking beam at 7:30 AM.

At 9:18 AM, Operations halted stacking while they investigated a problem with an Accumulator RF station (ARF1). Operators contacted an expert. Stacking resumed at 9:56 AM.

At 10:07 AM, Operations reported that the I- Source’s output had been drifting down for a while. An operator increased the extractor voltage and completed a 750 KeV Line tune followed by at Booster injection tune.

At 12:35 PM, a NuMI experimenter reported a LCW problem with its Horn. Operations left a message with a Horn expert. At 1:24 PM, Operations turned off beam to NuMI.

The Recycler began preparing to stash at 2:03 PM. The stash ended. Pbar resumed stacking and MCenter resumed taking beam around 3:20 PM.

Operations terminated store 4638 at 5:02 PM and put the TeV into shot setup a few minutes later.

Operations resumed sending beam to NuMI at 5:21 PM. Operations turned off NuMI at 5:51 PM for a NuMI access.

Operations established store 4639 at 7:14 PM with an initial luminosity of 166.91E30. Pbar resumed stacking and MCenter resumed taking beam.

Operations resumed sending beam to NuMI at 8:49 PM.

The Recycler began preparing to stash at 9:10 PM. The stash ended at 9:49 PM. Pbar resumed stacking and MCenter resumed taking beam at 10:01 PM.

At 11:47 PM, I- Source problems returned. Operations called in an expert.

Monday February 13

The midnight shift began with Operations monitoring store 4639, the stash and stack, and with all other beam down due to a problem with the Preaccelerator.

A Linac expert arrived and switching over to the H- Source at 1:01 AM.

The Proton Source resumed sending beam to Pbar, NuMI, MiniBooNE, and MCenter at 1:17 AM.

The Recycler began preparing to stash at 3:12 AM. The stash ended at 3:25AM. Pbar resumed stacking and MCenter resumed taking beam around 3:37 AM.

The Recycler began preparing to stash at 7:07 AM. The stash ended at 7:35 AM. Pbar resumed stacking and MCenter resumed taking beam around 7:49 AM.

The Plans for Monday

The plans for today are to stack, stash, and store.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/13/2006    email Fermilab